OpenStudy (anonymous):

HELP ALGEBRA 2 WILL MEDAL AND FAN!!!!!!!!!!!!!!

OpenStudy (anonymous):

\[(16x ^{8})^{-0.75}\] simplify

OpenStudy (nathan917):

1. (16x^8)^-0.75 = (2^4x^8)^(-3/4) Taking the 4th root you get: = (2x^2)^-3 Cubing you get: = (8x^6)^(-1) Invert to get: = 1/(8x^6)